# Khatri-Rao Clustering for Data Summarization

### v1

@@ 1. Introduction
Distilling datasets into succinct and meaningful summaries represents an essential and ubiquitous task. Clustering has emerged as a core methodology for tackling this problem. In particular, centroid-based clustering holds a central position when looking for effective data-summarization strategies. Given a notion of similarity, centroid-based clustering algorithms summarize data by assigning each data point to the closest element in a succinct set of representative data points called centroids. The data summaries provided by centroid-based clustering identify meaningful clusters and reveal patterns useful for applications such as exploratory data analysis, segmentation, anomaly detection, and document organization (Xu and Tian, 2015).
@@
In many cases, the interest lies in extracting a data summary which is as succinct and accurate as possible. Clustering has consistently shown success in summarizing data across diverse domains, for example, in image (Omari et al., 2025), document (Dvorskỳ et al., 2010), tabular (Han and Li, 2016), and spatio-temporal (Whelan et al., 2010) data processing, as well as for compressing deep neural networks (Son et al., 2018; Somasundaram and Rani, 2011).
@@
Pushing the boundaries of centroid-based clustering for data summarization through Khatri-Rao operators. Many modern applications demand the summarization of datasets of unprecedented scale (Draganov et al., 2024). As the size of a dataset increases, the underlying cluster structure may evolve either by enlarging clusters or by introducing additional clusters. Thus, modern datasets can exhibit a massive number of underlying clusters (Kobren et al., 2017; Bateni et al., 2021). For instance, studies on protein structures analyze millions of clusters (Barrio-Hernandez et al., 2023). Similarly, a large number of clusters can be incurred in topic modeling for documents (Sethia et al., 2022). In entity resolution, including record linkage and de-duplication, the number of clusters grows with dataset size (Betancourt et al., 2016), and in real-world networks the number of clusters tends to follow a power-law distribution that also increases with network size (Clauset et al., 2004). Centroid-based clustering struggles to yield data summaries that are both accurate and succinct in similar large-scale applications characterized by a massive number of clusters.
@@
In spite of this, no existing work challenges the long-standing centroid-based clustering paradigm toward more succinct data summarization. To address this gap, we adopt a data-compression perspective on centroid-based clustering, and we investigate the following research question: do standard centroid-based clustering algorithms produce data summaries that carry some degree of redundancy, suggesting the potential for further compression? To shed light on this question, we relax a crucial assumption of centroid-based clustering, namely that centroids are independent entities. Often, richer representations emerge from the interaction of simpler building blocks, as illustrated, e.g., by multi-head attention (Vaswani et al., 2017). In this spirit, we postulate that centroids admit a more succinct representation in terms of a smaller set of building blocks. More specifically, we introduce the Khatri-Rao clustering paradigm, in which centroids arise from the interaction of succinct sets of protocentroids through so-called Khatri-Rao operators. These operators aggregate each protocentroid in each set with all protocentroids of the other sets via elementwise operations such as sums or products.
@@
A simple example of this formulation is given in Figure 1. It shows that the nine clusters in the stickfigures dataset (Günnemann et al., 2014) can be represented by two sets of three protocentroids. Three protocentroids explain the upper part of the stick figures and three explain the lower part. The final centroids can then be generated by additively combining each protocentroid in one set with every protocentroid in the other set. In this case, centroids are said to exhibit a Khatri-Rao structure, and more precisely they correspond to the Khatri-Rao sum of the two sets of protocentroids. The stick-figure example shows that the dataset can be summarized by just $6$ images, whereas standard centroid-based clustering algorithms require $9$ images. Extending beyond the example, $p$ sets of $h_{1},h_{2},\dots$ and $h_{p}$ protocentroids can represent up to $\prod_{i=1}^{p}h_{i}$ centroids.
@@
As a first concrete instantiation of the Khatri-Rao clustering paradigm, we design Khatri-Rao-$k$-Means, which builds on the most popular algorithm for centroid-based clustering, i.e., the $k$-Means algorithm (Lloyd, 1982), to find centroids that can be succinctly expressed as the Khatri-Rao sum or product of two or more sets of protocentroids. Although Khatri-Rao-$k$-Means can result in considerably more succinct and yet equally accurate data summaries than standard $k$-Means, it is more prone to converge to undesirable local minima. Finding high-quality solutions may require exploring many different initializations. Motivated by this limitation, we introduce the Khatri-Rao deep clustering framework, which starts from Khatri-Rao-$k$-Means and extends the Khatri-Rao paradigm to deep clustering. Deep clustering effectively handles large-scale and high-dimensional data by incorporating deep-learning-based representation learning. By aligning the learned representations with the Khatri–Rao structure, the Khatri-Rao deep clustering framework overcomes the limitations of Khatri-Rao-$k$-Means, and usually provides data summaries that have a comparable accuracy to unconstrained baselines but are consistently more succinct.
Figure 2 anticipates the benefits of Khatri-Rao clustering by comparing $k$-Means and the deep clustering algorithms Improved Deep Embedded Clustering (IDEC) (Guo et al., 2017) and Deep-k-Means (DKM) (Fard et al., 2020) against their extensions based on the Khatri-Rao paradigm. The Khatri-Rao clustering algorithms maintain comparable accuracy (here measured by unsupervised clustering accuracy (Yang et al., 2010)) and simultaneously achieve a high level of compression.
@@ 2. Related Work
Clustering, data summarization and compression. Obtaining data summaries that are as succinct and as accurate as possible has established itself as a central problem in such domains as database systems under various names. It is useful to distinguish between data summarization and compression. The goal of data summarization is to craft a summary of a dataset that captures its essential patterns. Centroid-based clustering is a popular approach for data summarization, although alternative approaches exist (e.g., aggregation, dimensionality reduction, or sampling) and may better suit particular applications (Ahmed, 2019). A related field to data summarization is data compression, which emphasizes reducing data size for storage or transmission, either with or without information loss, over capturing the main patterns in the data (Jayasankar et al., 2021). The Khatri-Rao clustering paradigm effectively compresses the data summaries given by centroid-based clustering. Thus, our work pertains to both data summarization and compression. We henceforth use “compression” to refer specifically to the reduction of the size of a data summary.
@@
Research has leveraged information-theoretic principles to establish a theoretical foundation for understanding clustering as a compression problem, explicitly considering the trade-off between compression, efficiency and clustering quality (Bischof et al., 1999; Slonim et al., 2005; Raj and Wiggins, 2009). However, there is a lack of practical methods designed to enhance the compression capabilities of clustering, a gap we aim to address in this work. We concentrate on two central approaches to centroid-based clustering, $k$-Means and deep clustering.
@@
$k$-Means clustering. $k$-Means is a seminal problem in clustering (Lloyd, 1982), and the classic $k$-Means algorithm has established itself as the cornerstone of many clustering algorithms (Jain, 2008). More details on the $k$-Means problem and on the classic $k$-Means algorithm are given in Section 3. Today, $k$-Means clustering is still an active topic of research (Ikotun et al., 2023). Nonetheless, prior to our work which extends $k$-Means to the Khatri-Rao paradigm, improving the compression power of $k$-Means has been largely overlooked.
@@
Deep clustering. Unlike $k$-Means, which performs clustering directly in the input space, recent deep clustering algorithms combine clustering with neural networks to learn representations more amenable to clustering (Leiber et al., 2023; Ren et al., 2024; Zhou et al., 2025).
@@
Deep clustering algorithms can use various types of neural networks. We focus on autoencoder-based algorithms as they are agnostic to data type and easily extendable (Leiber et al., 2025). Centroid-based deep clustering algorithms are useful for summarization purposes and have demonstrated state-of-the-art performance (Miklautz et al., 2025). A well-known representative is Deep-k-Means (DKM) (Fard et al., 2020). DKM is similar to $k$-Means. However, it introduces soft cluster assignments based on the softmax function. Deep Embedded Clustering (DEC) (Xie et al., 2016) and its successor Improved Deep Embedded Clustering (IDEC) (Guo et al., 2017) utilize the Kullback-Leibler divergence (Kullback and Leibler, 1951) to learn an embedding that aligns a Student’s t model of the data distribution with a target distribution. Our work extends the DKM and IDEC algorithms to the Khatri-Rao paradigm. Further details on deep clustering, DKM and IDEC are given in Section 3.
@@
Matrix decomposition. Clustering is closely related to matrix decomposition, which expresses a matrix as a product of two simpler matrices. It is known that, given a data matrix, the model imposed by a centroid-based clustering algorithm like $k$-Means can be seen as a special case of a matrix-decomposition model where one of the factor matrices is constrained to indicate cluster assignments and the other stores centroids (Ding et al., 2005). Similarly, Khatri-Rao clustering hinges on known operators like Hadamard (i.e., elementwise) and Khatri-Rao products (Khatri and Rao, 1968) and is closely connected to the Hadamard decomposition, which models a data matrix as the Hadamard product of matrix decompositions (Ciaperoni et al., 2024). In particular, the model imposed by Khatri-Rao clustering with protocentroids aggregated via elementwise product on a data matrix is equivalent to a Hadamard decomposition where one matrix in each decomposition is constrained to indicate the assignment to protocentroids. In general, algorithms for matrix decomposition can be useful for clustering (Li and Ding, 2013) and can also be potentially extended to the Khatri-Rao clustering setting.
@@ 3. Preliminaries
We consider datasets $\mathcal{D}\subset\mathbb{R}^{m}$ composed of a set of data points $\{\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{n}\}$, where $\mathbf{x}_{i}=(x_{1},x_{2}\dots,x_{m})$. The goal of clustering is to divide the dataset into $k$ clusters $\{C_{1},C_{2},\ldots,C_{k}\}$ so that $\mathcal{D}=\cup_{i=1}^{k}C_{i}$ and $C_{i}\cap_{i\neq j}C_{j}=\emptyset$. Furthermore, we denote the set of cluster centroids by $M_{\boldsymbol{\mu}}=\{\boldsymbol{\mu}_{1},\boldsymbol{\mu}_{2},\dots,\boldsymbol{\mu}_{k}\}\subset\mathbb{R}^{m}$. In the Khatri-Rao clustering paradigm, we define the centroids in $M_{\boldsymbol{\mu}}$ more succinctly by combining protocentroids. Unless stated otherwise, we assume that there are $k$ centroids and $p$ sets of protocentroids, where the $i$-th set has cardinality $h_{i}$. The protocentroids in different sets are combined using a function referred to as aggregator and denoted by $\oplus$. While the Khatri-Rao clustering paradigm is general and could in principle accommodate any aggregator function, in this work, we focus on the sum and product, i.e., $\oplus\in\{+,\times\}$. When $\oplus$ is left unspecified, it is understood to represent an arbitrary choice of aggregator. For notational convenience, $\oplus$ is applied to scalars, vectors, matrices and sets. When applied to vectors and matrices, it is an elementwise operator, corresponding to the standard sum for $\oplus=+$ and to the Hadamard product for $\oplus=\times$. When applied to sets (e.g., sets of protocentroids), it is a so-called Khatri-Rao operator. We define the Khatri-Rao $\oplus$ operator as an operator that, given $p$ sets of vectors, produces the set of all vectors obtained by elementwise application of $\oplus$ to all possible combinations with one vector from each set. The name “Khatri-Rao” operator is chosen since, if protocentroids in each set are arranged as rows of matrices and $\oplus=\times$, the operator reduces to the Khatri-Rao product (Khatri and Rao, 1968).
@@
$k$-Means clustering. In $k$-Means, the dataset is summarized in terms of a set of centroids $\Theta=\{\boldsymbol{\mu}_{i}\}_{i=1}^{k}$, and the measure of quality is the total squared Euclidean distance of each point to the closest centroid, henceforth called inertia, as is standard in the literature (Pedregosa et al., 2011).
@@
To address this problem, the standard $k$-Means algorithm starts by initializing $k$ cluster centroids. Centroids can be initialized by sampling data points uniformly at random. Alternatively, the popular $k\textsc{-Means++}$ (Arthur and Vassilvitskii, 2007) strategy chooses data points that are sufficiently far apart from each other as initial centroids, which gives theoretical approximation guarantees and often results in performance improvements. After initialization, the $k$-Means algorithm iteratively assigns each data point to its nearest (in Euclidean distance) centroid and updates the cluster centroids by computing the mean of the points assigned to each cluster. The algorithm terminates and outputs the current centroids when either the centroids converge to a stable position or a maximum number of iterations is reached.
@@
where $\Theta_{\boldsymbol{\mu}}=\{\boldsymbol{\mu}_{1},\dots,\boldsymbol{\mu}_{k}\}$ denotes the cluster centroids in the latent space and $w_{rec}\geq 0$ balances clustering and representation learning. In our study, we consider two clustering loss functions for $\mathcal{L}_{\mathrm{cluster}}$; the ones proposed for the DKM (Fard et al., 2020) and IDEC (Guo et al., 2017) algorithms. Thus, $\mathcal{L}_{\mathrm{cluster}}\in\{\mathcal{L}_{\mathrm{DKM}},\mathcal{L}_{\mathrm{IDEC}}\}$, where
@@
with $q_{l,i}$ and $p_{l,i}$ representing the data and target distributions, respectively, and $a$ is an algorithm-specific parameter that is usually set to $1$ for IDEC and $1000$ for DKM. To obtain a succinct data summary in terms of autoencoder and centroid parameters, the objective in Equation (2) is typically optimized via batch-wise backpropagation, using automatic differentiation (Leiber et al., 2023).

### v2

@@ 1. Introduction
Distilling datasets into succinct and meaningful summaries represents an essential and ubiquitous task. Clustering has emerged as a core methodology for tackling this problem. In particular, centroid-based clustering holds a central position when looking for effective data-summarization strategies. Given a notion of similarity, centroid-based clustering algorithms summarize data by assigning each data point to the closest element in a succinct set of representative data points called centroids. The data summaries provided by centroid-based clustering identify meaningful clusters and reveal patterns useful for applications such as exploratory data analysis, segmentation, anomaly detection, and document organization (xu2015comprehensive).
@@
In many cases, the interest lies in extracting a data summary which is as succinct and accurate as possible. Clustering has consistently shown success in summarizing data across diverse domains, for example, in image (omari2025advancing), document (dvorsky2010document), tabular (han2016lossless), and spatio-temporal (whelan2010data) data processing, as well as for compressing deep neural networks (son2018clustering; somasundaram2011novel).
@@
Pushing the boundaries of centroid-based clustering for data summarization through Khatri-Rao operators. Many modern applications demand the summarization of datasets of unprecedented scale (draganov2024settling). As the size of a dataset increases, the underlying cluster structure may evolve either by enlarging clusters or by introducing additional clusters. Thus, modern datasets can exhibit a massive number of underlying clusters (kobren2017hierarchical; bateni2021extreme). For instance, studies on protein structures analyze millions of clusters (barrio2023clustering). Similarly, a large number of clusters can be incurred in topic modeling for documents (sethia2022framework). In entity resolution, including record linkage and de-duplication, the number of clusters grows with dataset size (betancourt2016flexible), and in real-world networks the number of clusters tends to follow a power-law distribution that also increases with network size (clauset2004finding). Centroid-based clustering struggles to yield data summaries that are both accurate and succinct in similar large-scale applications characterized by a massive number of clusters.
@@
In spite of this, no existing work challenges the long-standing centroid-based clustering paradigm toward more succinct data summarization. To address this gap, we adopt a data-compression perspective on centroid-based clustering, and we investigate the following research question: do standard centroid-based clustering algorithms produce data summaries that carry some degree of redundancy, suggesting the potential for further compression? To shed light on this question, we relax a crucial assumption of centroid-based clustering, namely that centroids are independent entities. Often, richer representations emerge from the interaction of simpler building blocks, as illustrated, e.g., by multi-head attention (vaswani2017attention). In this spirit, we postulate that centroids admit a more succinct representation in terms of a smaller set of building blocks. More specifically, we introduce the Khatri-Rao clustering paradigm, in which centroids arise from the interaction of succinct sets of protocentroids through so-called Khatri-Rao operators. These operators aggregate each protocentroid in each set with all protocentroids of the other sets via elementwise operations such as sums or products.
@@
A simple example of this formulation is given in Figure 1. It shows that the nine clusters in the stickfigures dataset (gunnemann2014smvc) can be represented by two sets of three protocentroids. Three protocentroids explain the upper part of the stick figures and three explain the lower part. The final centroids can then be generated by additively combining each protocentroid in one set with every protocentroid in the other set. In this case, centroids are said to exhibit a Khatri-Rao structure, and more precisely they correspond to the Khatri-Rao sum of the two sets of protocentroids. The stick-figure example shows that the dataset can be summarized by just $6$ images, whereas standard centroid-based clustering algorithms require $9$ images. Extending beyond the example, $p$ sets of $h_{1},h_{2},\dots$ and $h_{p}$ protocentroids can represent up to $\prod_{i=1}^{p}h_{i}$ centroids.
@@
As a first concrete instantiation of the Khatri-Rao clustering paradigm, we design Khatri-Rao-$k$-Means, which builds on the most popular algorithm for centroid-based clustering, i.e., the $k$-Means algorithm (lloyd1982least), to find centroids that can be succinctly expressed as the Khatri-Rao sum or product of two or more sets of protocentroids. Although Khatri-Rao-$k$-Means can result in considerably more succinct and yet equally accurate data summaries than standard $k$-Means, it is more prone to converge to undesirable local minima. Finding high-quality solutions may require exploring many different initializations. Motivated by this limitation, we introduce the Khatri-Rao deep clustering framework, which starts from Khatri-Rao-$k$-Means and extends the Khatri-Rao paradigm to deep clustering. Deep clustering effectively handles large-scale and high-dimensional data by incorporating deep-learning-based representation learning. By aligning the learned representations with the Khatri–Rao structure, the Khatri-Rao deep clustering framework overcomes the limitations of Khatri-Rao-$k$-Means, and usually provides data summaries that have a comparable accuracy to unconstrained baselines but are consistently more succinct.
Figure 2 anticipates the benefits of Khatri-Rao clustering by comparing $k$-Means and the deep clustering algorithms Improved Deep Embedded Clustering (IDEC) (Guo2017ImprovedDE) and Deep-k-Means (DKM) (fard2020deep) against their extensions based on the Khatri-Rao paradigm. The Khatri-Rao clustering algorithms maintain comparable accuracy (here measured by unsupervised clustering accuracy (Yang2010accuracy)) and simultaneously achieve a high level of compression.
@@ 2. Related Work
Clustering, data summarization and compression. Obtaining data summaries that are as succinct and as accurate as possible has established itself as a central problem in such domains as database systems under various names. It is useful to distinguish between data summarization and compression. The goal of data summarization is to craft a summary of a dataset that captures its essential patterns. Centroid-based clustering is a popular approach for data summarization, although alternative approaches exist (e.g., aggregation, dimensionality reduction, or sampling) and may better suit particular applications (ahmed2019data). A related field to data summarization is data compression, which emphasizes reducing data size for storage or transmission, either with or without information loss, over capturing the main patterns in the data (jayasankar2021survey). The Khatri-Rao clustering paradigm effectively compresses the data summaries given by centroid-based clustering. Thus, our work pertains to both data summarization and compression. We henceforth use “compression” to refer specifically to the reduction of the size of a data summary.
@@
Research has leveraged information-theoretic principles to establish a theoretical foundation for understanding clustering as a compression problem, explicitly considering the trade-off between compression, efficiency and clustering quality (bischof1999mdl; slonim2005information; raj2009information). However, there is a lack of practical methods designed to enhance the compression capabilities of clustering, a gap we aim to address in this work. We concentrate on two central approaches to centroid-based clustering, $k$-Means and deep clustering.
@@
$k$-Means clustering. $k$-Means is a seminal problem in clustering (lloyd1982least), and the classic $k$-Means algorithm has established itself as the cornerstone of many clustering algorithms (jain2008data). More details on the $k$-Means problem and on the classic $k$-Means algorithm are given in Section 3. Today, $k$-Means clustering is still an active topic of research (ikotun2023k). Nonetheless, prior to our work which extends $k$-Means to the Khatri-Rao paradigm, improving the compression power of $k$-Means has been largely overlooked.
@@
Deep clustering. Unlike $k$-Means, which performs clustering directly in the input space, recent deep clustering algorithms combine clustering with neural networks to learn representations more amenable to clustering (leiber2023benchmarking; ren2024deep; zhou2025comprehensive).
@@
Deep clustering algorithms can use various types of neural networks. We focus on autoencoder-based algorithms as they are agnostic to data type and easily extendable (leiber2025introductory). Centroid-based deep clustering algorithms are useful for summarization purposes and have demonstrated state-of-the-art performance (miklautz2025breaking). A well-known representative is Deep-k-Means (DKM) (fard2020deep). DKM is similar to $k$-Means. However, it introduces soft cluster assignments based on the softmax function. Deep Embedded Clustering (DEC) (xie2016unsupervised) and its successor Improved Deep Embedded Clustering (IDEC) (Guo2017ImprovedDE) utilize the Kullback-Leibler divergence (kullback1951information) to learn an embedding that aligns a Student’s t model of the data distribution with a target distribution. Our work extends the DKM and IDEC algorithms to the Khatri-Rao paradigm. Further details on deep clustering, DKM and IDEC are given in Section 3.
@@
Matrix decomposition. Clustering is closely related to matrix decomposition, which expresses a matrix as a product of two simpler matrices. It is known that, given a data matrix, the model imposed by a centroid-based clustering algorithm like $k$-Means can be seen as a special case of a matrix-decomposition model where one of the factor matrices is constrained to indicate cluster assignments and the other stores centroids (ding2005equivalence). Similarly, Khatri-Rao clustering hinges on known operators like Hadamard (i.e., elementwise) and Khatri-Rao products (khatri1968solutions) and is closely connected to the Hadamard decomposition, which models a data matrix as the Hadamard product of matrix decompositions (ciaperoni2024hadamard). In particular, the model imposed by Khatri-Rao clustering with protocentroids aggregated via elementwise product on a data matrix is equivalent to a Hadamard decomposition where one matrix in each decomposition is constrained to indicate the assignment to protocentroids. In general, algorithms for matrix decomposition can be useful for clustering (li2013nonnegative) and can also be potentially extended to the Khatri-Rao clustering setting.
@@ 3. Preliminaries
We consider datasets $\mathcal{D}\subset\mathbb{R}^{m}$ composed of a set of data points $\{\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{n}\}$, where $\mathbf{x}_{i}=(x_{1},x_{2}\dots,x_{m})$. The goal of clustering is to divide the dataset into $k$ clusters $\{C_{1},C_{2},\ldots,C_{k}\}$ so that $\mathcal{D}=\cup_{i=1}^{k}C_{i}$ and $C_{i}\cap_{i\neq j}C_{j}=\emptyset$. Furthermore, we denote the set of cluster centroids by $M_{\boldsymbol{\mu}}=\{\boldsymbol{\mu}_{1},\boldsymbol{\mu}_{2},\dots,\boldsymbol{\mu}_{k}\}\subset\mathbb{R}^{m}$. In the Khatri-Rao clustering paradigm, we define the centroids in $M_{\boldsymbol{\mu}}$ more succinctly by combining protocentroids. Unless stated otherwise, we assume that there are $k$ centroids and $p$ sets of protocentroids, where the $i$-th set has cardinality $h_{i}$. The protocentroids in different sets are combined using a function referred to as aggregator and denoted by $\oplus$. While the Khatri-Rao clustering paradigm is general and could in principle accommodate any aggregator function, in this work, we focus on the sum and product, i.e., $\oplus\in\{+,\times\}$. When $\oplus$ is left unspecified, it is understood to represent an arbitrary choice of aggregator. For notational convenience, $\oplus$ is applied to scalars, vectors, matrices and sets. When applied to vectors and matrices, it is an elementwise operator, corresponding to the standard sum for $\oplus=+$ and to the Hadamard product for $\oplus=\times$. When applied to sets (e.g., sets of protocentroids), it is a so-called Khatri-Rao operator. We define the Khatri-Rao $\oplus$ operator as an operator that, given $p$ sets of vectors, produces the set of all vectors obtained by elementwise application of $\oplus$ to all possible combinations with one vector from each set. The name “Khatri-Rao” operator is chosen since, if protocentroids in each set are arranged as rows of matrices and $\oplus=\times$, the operator reduces to the Khatri-Rao product (khatri1968solutions).
@@
$k$-Means clustering. In $k$-Means, the dataset is summarized in terms of a set of centroids $\Theta=\{\boldsymbol{\mu}_{i}\}_{i=1}^{k}$, and the measure of quality is the total squared Euclidean distance of each point to the closest centroid, henceforth called inertia, as is standard in the literature (scikit-learn).
@@
To address this problem, the standard $k$-Means algorithm starts by initializing $k$ cluster centroids. Centroids can be initialized by sampling data points uniformly at random. Alternatively, the popular $k\textsc{-Means++}$ (arthur2007k) strategy chooses data points that are sufficiently far apart from each other as initial centroids, which gives theoretical approximation guarantees and often results in performance improvements. After initialization, the $k$-Means algorithm iteratively assigns each data point to its nearest (in Euclidean distance) centroid and updates the cluster centroids by computing the mean of the points assigned to each cluster. The algorithm terminates and outputs the current centroids when either the centroids converge to a stable position or a maximum number of iterations is reached.
@@
where $\Theta_{\boldsymbol{\mu}}=\{\boldsymbol{\mu}_{1},\dots,\boldsymbol{\mu}_{k}\}$ denotes the cluster centroids in the latent space and $w_{rec}\geq 0$ balances clustering and representation learning. In our study, we consider two clustering loss functions for $\mathcal{L}_{\mathrm{cluster}}$; the ones proposed for the DKM (fard2020deep) and IDEC (Guo2017ImprovedDE) algorithms. Thus, $\mathcal{L}_{\mathrm{cluster}}\in\{\mathcal{L}_{\mathrm{DKM}},\mathcal{L}_{\mathrm{IDEC}}\}$, where
@@
with $q_{l,i}$ and $p_{l,i}$ representing the data and target distributions, respectively, and $a$ is an algorithm-specific parameter that is usually set to $1$ for IDEC and $1000$ for DKM. To obtain a succinct data summary in terms of autoencoder and centroid parameters, the objective in Equation (2) is typically optimized via batch-wise backpropagation, using automatic differentiation (leiber2023benchmarking).